OpenStudy (anonymous):

Solve the equation 1 over 2 x + 2 = 22 for x.

OpenStudy (anonymous):

1/2x + 2 = 22 1/2x + 2 - 2 = 22 - 2 1/2x = 20 1/2x * 2 = 20 * 2 x = 40

OpenStudy (anonymous):

thank you!! (: Solve the equation 2 over 5 (x − 2) = 4x for x. @FriedRice

OpenStudy (anonymous):

well its the same thing 2/5(x - 2) = 4x 2/5(x - 2)*5/2 = 4x * 5/2 x - 2 = 10x -2 = 9x x = -2/9